OpenStudy (anonymous):

A 200 g particle gains a momentum of 25 Ns in 12 s starting from rest. Find the kinetic energy of the particle and the applied force.help !

OpenStudy (anonymous):

Momentum is given by mass *velocity. Taking 25/ 0.2 =125m/s Kinetic energy = 0.5*m *v^2 0.5*0.2*125^2=1562.5J Applied force = mass * acceleration. since it took 12 seconds for the particle to go from 0 m/s to 125m/s, acceleration= 125/12= 10.4m/s^2 Force = 0.2 * 10.4 2.08N

OpenStudy (anonymous):

can u help my latest question tiaph ?